# Periodic Korteweg-de Vries soliton potentials generate magnetic field strength with excellent quasisymmetry

### v2

@@ III Perturbative solution to the overdetermination problem of QS in vacuum magnetic fields
Vacuum fields with a very high degree of QS, obtained through numerical optimization [4], appear to be characterized by exceptionally low magnetic shear. Since magnetic shear plays an essential role in MHD stability theory, it is important to understand if QS is inextricably tied to the low shear of such configurations.
The second-order NAE theory predictions match the precise QS configurations remarkably well [54, 12]. However, the second-order NAE theory cannot readily capture the effects of magnetic shear for quasisymmetric configurations. The reason behind this is connected to the overdetermination problem. While magnetic shear near the axis needs third and fourth-order contributions [23, 24, 25] in the NAE, the quasisymmetric system becomes hopelessly overdetermined beyond second-order.
We will demonstrate here that it is possible to address both the issues of overdetermination and weak magnetic shear. We begin with the observation that the remarkable accuracy with which the flux surface shapes and the magnetic field strength are captured within second-order NAE suggests a multi-scale approach to the QS problem, where the magnetic shear introduces another length scale. Let $\alpha=\theta-\iota(\psi)\phi$ denote the field line label in a straight field line coordinate system $(\psi,\theta,\phi)$. Its gradient, $\bm{\nabla}\alpha$, is multi-valued and given by
@@
Near the axis, $|\bm{\nabla}\theta|\sim\rho^{-1}$, where $\rho=\sqrt{2\psi}$ is the distance from the axis, whereas $|\bm{\nabla}\phi|\sim R^{-1}$, where $R$ is the typical radius of curvature of the axis and $\rho\ll R$ in the NAE. Thus, the length scale $L^{-1}_{\alpha 1}\sim\rho^{-1}$ near the axis, which represents the typical scale of poloidal variations of surfaces. On the other hand, $L^{-1}_{\alpha 2}\sim\iota^{\prime}(\psi)|\bm{\nabla}\psi|\sim\hat{s}\rho$ near the axis, where $\hat{s}$ is the on-axis shear. Thus, $L_{\alpha 2}$ is greater than $L_{\alpha 1}$ by (typically, approximately two) orders of magnitude in the distance from the axis. In particular, for second-order NAE, $L_{\alpha 2}$ tends to infinity since $\iota$ is approximated by the on-axis rotational transform $\iota_{0}$. As we move further away from the axis, a finite value of $L_{\alpha 2}$ is needed to capture the effect of shear. The two disparate scales $L_{\alpha 1},L_{\alpha 2}$ clearly provide a basis for a multi-scale approach, akin to the well-known approach in the theory of ballooning modes [55, 56, 57].
@@
We propose a free-surface approach to QS with the two scales in mind. We assume that the second-order NAE theory has been used to construct a good volumetric quasisymmetric vacuum field. The NAE solution serves as our inner solution near the axis. The outer free-surface solution arises from the surface expansion approach. Consistent with the observation that NAE predictions for the geometry are surprisingly accurate, we shall assume that the surface expansion can be carried out in a small region around a surface generated using the second-order NAE. The surface expansion allows us to bypass the difficulties of including the effects of shear within the framework of the NAE. The outermost flux surface will be treated as a free surface determined self-consistently by imposing QS and force balance.
@@
An essential step in any multi-scale analysis is to average over the fast scale to obtain an equation that involves only the slow scale. For example, in the study of ballooning modes near the axis [55], a poloidal averaging of the standard ballooning equation is carried out to average over the fast variation $L_{\alpha 1}\sim\rho$ and obtain a “distilled ballooning equation”, which involves only the longer length scale $L_{\alpha 2}\sim\hat{s}\rho$. A rigorous multiscale NAE program must be carried out for this purpose, which is beyond the scope of the present work. Here, we shall content ourselves with geometries that depend only weakly on the fast scale, thereby bypassing the need for averaging over the fast scale to obtain the “distilled” equations. Vacuum fields in slab and cylindrical geometries offer such possibilities.
@@
In Section III.1, we first demonstrate the free-surface approach to solving the overdetermination problem of QS in a vacuum field for slab geometry. The analysis is simplified, and lays bare the balance between dispersion and nonlinearity needed to realize soliton solutions [58]. Once this structure is understood, it essentially carries over with minimal modification to the more general geometry, which we discuss in III.2.
@@ III.1 Vacuum quasisymmetric magnetic field in slab geometry
where $\hat{q}$ is a small deviation from the low-order rational. The inner solution is trivially quasisymmetric. We will then carry out a surface expansion around the plane $z=h$ with the outermost free-surface parametrized by $z=h+\eta(x,y)$. The amplitude of $\eta(x,y)$ and $h$ will be assumed small compared to the characteristic length $L$ along the field line, which allows us to carry out a multiscale analysis for a weakly nonlinear system typical in the theory of shallow-water waves [58]. The justification and the relevance of the assumptions will be made clear in the following.
Exact vacuum fields in slab geometry with closed field lines and nested flux surfaces have been obtained in [59]. As shown in [59], one of the peculiarities of closed field line systems in a slab geometry is that one can always re-orient the coordinate system such that the magnetic field is purely along one of the axes. For a closed field line, tangential to the $X,Y$ plane and $\iota\equiv(\bm{B}\cdot\bm{\nabla}Y)/(\bm{B}\cdot\bm{\nabla}X)=n/m$, this can be achieved through a combination of rotation and dilatation of the coordinates
@@
The system (24) is mathematically similar to some aspects of the theory of shallow-water waves under gravity. Since our problem has no explicit time, we make $\alpha$ the analog of time. Therefore, in analogy with the shallow-water wave theory [58], we denote the condition (24d) as the “kinematic boundary condition” (KBC), and (24c) as the “generalized Bernoulli” (GB) condition. The differences between water wave theory and QS arise from the difference in the Bernoulli condition and the fact that we treat $y$ (approximately $\alpha$) as a time-like variable.
@@
Guided by the “principle of maximal balance” [60, 58], we order
@@
In the water-wave theory, the relation (26) represents the delicate balance between nonlinearity and dispersion, which is crucial for the existence of solitons. The significance of (26) in the context of vacuum fields can be clarified by NAE theory. The small parameter $\delta$ is analogous to the near-axis expansion parameter $\kappa\rho$, which is the product of the axis curvature $\kappa$ and distance $\rho$ from the axis. The balance (26) then implies that there are second-order corrections to the NAE flux surface shapes due to the magnetic shear. The difference between the actual surface and the second-order NAE predictions being $O(\delta^{2})$ is consistent with the results obtained using NAE and standard numerical optimization [54, 12]. We note that (26) is not the only possible relation for balancing nonlinearity and dispersion [61]. Other possibilities will be discussed in a future publication.
@@
Here, $Y=\epsilon y$ denotes a longer length scale in $y$. The TW solution is consistent with (6) expanded to the lowest order in $\epsilon$.
@@
Proceeding with the KBC (29d), we find that the lowest order
@@
Through suitable variable transformations, the above equation can be cast as the standard KdV equation for $B=1+\epsilon\tilde{B}$
@@
Let us point out some of the important predictions of the slab theory. Firstly, the sum $B_{M}+B_{m}+B_{X}$ is determined by the coefficients of KdV [30], which, as seen from (40) only depends on $q_{0}$. Moreover, as predicted by the NAE theory (Section V) and confirmed numerically (Section VI and Appendix E), the sum of $B_{M},B_{m}$ is essentially a constant equal to the field strength on the axis. Therefore, the third root $B_{X}$, must be determined by the $\iota$ profile. We shall also numerically verify this observation in Section VI.
@@
As shown in Appendix B subsidiary expansion leads to the Gardner’s equation [62, 33]
@@
where $c_{1}$ is a constant. Due to the additional $B^{2}$ term in the Gardner’s equation (46) compared to the KdV equation (41), under the TW ansatz we need to replace the cubic $f^{2}$ in (42) by a quartic [33]. Therefore, we need four roots to describe $B$ in this case, either all real, or two real and a pair of complex conjugates.
@@ III.2 Extension of the slab geometry results to general toroidal geometry
To include the effect of shear, we now go beyond the NAE limit and extend the free-surface idea proposed earlier in the slab model to more general geometries. We seek geometries in which the dependence on the fast variation is straightforward. Such a case arises with a large-aspect-ratio torus with nearly circular flux surfaces. Instead of the Cartesian coordinates, we shall now use the Clebsch coordinates $(\psi_{s},\alpha_{s},\Phi_{s})$, where the subscript denotes the surface around which we carry out the surface expansion. In these NAE Clebsch coordinates, we have
@@
In summary, the perturbative approach to solving the overdetermined vacuum quasisymmetric system leads to the following results. First, we can show analytically that the compatibility condition between force balance and QS determines the rotational transform profile. Within our approximate model, we are able to show that the compatibility implies that for a given average rotational transform, there is a preference for an optimal shear. In fact, recent work [36] has shown that stellarators with precise levels of QS that can be derived from the Landreman-Paul precise QA indeed have an optimal shear for the best QS. We discuss this further in Section VI (see Fig. 5). Second, the functional form of $(\partial_{\ell}B)^{2}$ is not arbitrary. In general, it is cubic in $B$, which leads to the KdV equation for $B$. For small $\iota$ and $N=0$, $(\partial_{\ell}B)^{2}$ is a quartic, leading to Gardner’s equation for $B$.
@@ IV The connection of QS to integrable PDEs: Painlevé property
We note that the first assumption is physically reasonable, has been shown to hold near the axis [17], and is also borne out by extensive computations. The second assumption is an essential requirement for QS. Both of these assumptions do not apply to omnigeneity. Regarding the third assumption, an alternative is to demand that $B$ must satisfy the fundamental TW condition (6) everywhere on the flux surface. The third condition leads to an infinite number of field-line $\alpha$-independent integrals of the form (17). In the perturbative approach, in addition to these assumptions, we also made certain assumptions on the geometry to make analytical progress with the force balance condition.
@@
Before we proceed with our analysis, let us discuss some of the important known exceptional cases of QS. As stated earlier, the case with an exact symmetry (axisymmetric or helically symmetric) is a special case of QS because $\bm{u}$ is determined independently of $\bm{B}$. The third condition, or equivalently $\bm{u}\cdot\bm{\nabla}B=0$, is satisfied identically for every choice of $f(B,\psi)$ in the $\partial_{\ell}B$ equation (10). We find a similar degeneracy in approximate QS near axisymmetry [63, 19, 64], where no constraints exist on $\partial_{\ell}B$. The near-axisymmetric devices can not be made quasisymmetric to all orders and have higher symmetry-breaking errors than the Landreman-Paul precise QA. A more general analysis shows that in configurations that are not close to either axisymmetry or helical symmetry, the degeneracy is avoided, and $\partial_{\ell}B$ is once again constrained [65]. We shall also exclude the exceptional case of zero rotational transform [66] since we focus only on irrational surfaces while assuming continuity on rational surfaces. We shall return to the comparison with the axisymmetric case at the end of this Section.
@@
Our main goal is to study the single-valuedness of $B$ on a constant flux surface $\psi$ that satisfies the $\partial_{\ell}B$ equation (10) and the $\partial_{\alpha}B$ equation (12), subject to the periodicity condition (11). Starting with the pioneering work by S. Kovalevskaya [67] on rigid body dynamics, the study of single-valuedness and singularities of differential equations in the complex domain has yielded fundamental results in several physics problems from quantum field theory to optics [68]. Extending from the real to the complex plane is a helpful technique that physicists often utilize in solving linear problems. However, as Kovaleskaya showed, going to the complex plane and analyzing singularities of highly nonlinear systems can be equally insightful. Moreover, singularities in the complex plane determine and control the behavior of analytic functions in the real line. An important example of this is a doubly periodic elliptic function [69, 70] whose singularities determine its periodicity on the real line. Our investigation is primarily motivated by the observation that the hidden symmetry of QS could admit a special Kovalevskaya-like solution [40].
@@
To proceed further, we need critical ideas from the well-developed theory of complex ODEs [71, 69, 67]. In contrast to solutions of linear ODEs, whose singularities are determined by those of the coefficients, nonlinear ODEs allow movable singularities, which depend on the initial conditions. The standard example is $y^{\prime}+y^{2}=0$, whose general solution, $y=(x-x_{0})^{-1}$, depends on the integration constant $x_{0}$ determined by the initial condition $y(0)=-x_{0}^{-1}$. The possible singularities of ODEs are known. They include poles, branch points, essential singular points, and essential singular lines. Single-valuedness of functions breaks down near critical singularities, such as branch points of algebraic or logarithmic nature. If the critical singularities are fixed, the functions can be made single-valued by defining suitable branch cuts or fixed Riemann surfaces. However, the presence of singularities, which are both critical and movable, leads to dense multi-valuedness around movable singularities of solutions and nonintegrability ([67, 72] and references therein). The absence of movable critical singularities in the general solution of an ODE is called the Painlevé property (PP) of an ODE.
We now recall some classical results on ODEs with the PP (details given in [71, 69, 67]). Let us rewrite the $\partial_{\ell}B$ equation in the complex plane as
@@
which is exactly solvable in terms of elliptic functions and their degenerate cases. As is well-known, elliptic functions are doubly-periodic, meromorphic functions. The dependence on the arbitrary initial condition or the TW frame thus leads to only movable poles and no movable critical singularities. Hence, the solution possesses the PP. These classical results follow from standard Painlevé test and extensions of Malmquist’s theorem [69, 73]. They show how stringent and rigid the constraint of single-valuedness can be.
Returning to the QS problem, we first observe that the Riccati equation for $B$ is inconsistent with the NAE solution. The Riccati equation for $B$ will have constant real coefficients on a flux surface, which does not admit nontrivial periodic solutions on the real line [74]. As discussed in Section III.1 and III.2 the perturbative treatment near the slab or the axis shows that $(\partial_{\ell}B)^{2}=P_{3}(B),P_{4}(B)$ is more appropriate, which coincides with two of the four forms of ODEs of first order and second degree with PP. The periodicity condition can be easily satisfied with the elliptic functions for these cases. Now, the cubic and quartic polynomials are the twice-integrated forms of the KdV and Gardner’s equations. The PP of the KdV and Gardner’s equation in the TW frame is well-known and is conjectured to be essential for integrable PDEs with soliton solutions [75, 76].
We would now like to address whether PP is necessary and sufficient for a quasisymmetric $B$. It is certainly sufficient for QS since both the triple product form (8) and the periodicity condition can be satisfied. However, the necessity is not apparent, particularly in light of axisymmetry, where no such stringent restriction on $\partial_{\ell}B$ appears. In axisymmetry, weak singularities such as $R^{m}\log{R}^{n}$, with integer $m,n$ often arise in the description of $\psi$ [77] and in $B$ in cylindrical coordinates. However, these singularities are fixed and, therefore, not of the critical movable type. To further understand the difference between a perfectly axisymmetric system and a system with excellent but not exact QS, we need to look deeper into the $\bm{u}\cdot\bm{\nabla}B=0$ condition. As discussed in Appendix A, if QS were exact, we would have the same characteristics as axisymmetry. However, unlike the axisymmetric problem, the QS problem is intrinsically overdetermined and three-dimensional. With this in mind, we shall now motivate the PP as a criterion that allows one to look for robust quasisymmetric systems.
@@
Figure 1: The value of $(dB/d\ell)^{2}$ on the outermost flux surface in the precise QA equilibrium [4] is plotted as a function of $B$, alongside a cubic polynomial fit in (a). In (b), the errors of quadratic, cubic, and quartic fits are compared on each flux surface, demonstrating that a cubic ansatz (as predicted by KdV) is sufficient. Hence, precise QA has the Painlevé property for $B$.

### v3

@@ III Perturbative solution to the overdetermination problem of QS for vacuum magnetic fields
Vacuum fields with a very high degree of QS, obtained through numerical optimization Landreman and Paul (2022), appear to be characterized by exceptionally low magnetic shear. Since magnetic shear plays an essential role in MHD stability theory, it is important to understand if QS is inextricably tied to the low shear of such configurations.
The second-order NAE theory predictions match the precise QS configurations remarkably well (Landreman, 2022; Rodriguez, 2022). However, the second-order NAE theory cannot readily capture the effects of magnetic shear for quasisymmetric configurations. The reason behind this is connected to the overdetermination problem. While magnetic shear near the axis needs third and fourth-order contributions Rodríguez et al. (2022); Mercier (1974); Solov’ev and Shafranov (1970) in the NAE, the quasisymmetric system becomes hopelessly overdetermined beyond second-order.
We begin with the observation that the remarkable accuracy with which the flux surface shapes and the magnetic field strength are captured within second-order NAE suggests a multi-scale approach to the QS problem, where the magnetic shear introduces another length scale. Let $\alpha=\theta-\iota(\psi)\phi$ denote the field line label in a straight field line coordinate system $(\psi,\theta,\phi)$. Its gradient, $\bm{\nabla}\alpha$, is multi-valued and given by
@@
Near the axis, $|\bm{\nabla}\theta|\sim\rho^{-1}$, where $\rho=\sqrt{2\psi}$ is the distance from the axis, whereas $|\bm{\nabla}\phi|\sim R^{-1}$, where $R$ is the typical radius of curvature of the axis and $\rho\ll R$ in the NAE. Thus, the length scale $L^{-1}_{\alpha 1}\sim\rho^{-1}$ near the axis, which represents the typical scale of poloidal variations of surfaces. On the other hand, $L^{-1}_{\alpha 2}\sim\iota^{\prime}(\psi)|\bm{\nabla}\psi|\sim\hat{s}\rho$ near the axis, where $\hat{s}$ is the on-axis shear. Thus, $L_{\alpha 2}$ is greater than $L_{\alpha 1}$ by (typically, approximately two) orders of magnitude in the distance from the axis. In particular, for second-order NAE, $L_{\alpha 2}$ tends to infinity since $\iota$ is approximated by the on-axis rotational transform $\iota_{0}$. As we move further away from the axis, a finite value of $L_{\alpha 2}$ is needed to capture the effect of shear. The two disparate scales $L_{\alpha 1},L_{\alpha 2}$ clearly provide a basis for a multi-scale approach, akin to the well-known approach in the theory of ballooning modes (Antonsen Jr et al., 1982; Coppi et al., 1980; Hazeltine and Meiss, 1985).
@@
We propose a free-surface approach to QS with the two scales in mind. We assume that the second-order NAE theory has been used to construct a good volumetric quasisymmetric vacuum field. The NAE solution serves as our inner solution near the axis. The outer free-surface solution arises from the surface expansion approach. Consistent with the observation that NAE predictions for the geometry are surprisingly accurate, we shall assume that the surface expansion can be carried out in a small region around a surface, generated using the second-order NAE. The surface expansion allows us to bypass the difficulties of including the effects of shear within the framework of the NAE. The outermost flux surface will be treated as a free surface, determined self-consistently by imposing QS and force balance.
@@
An essential step in any multi-scale analysis is to average over the fast scale to obtain an equation that involves only the slow scale. For example, in the study of ballooning modes near the axis (Antonsen Jr et al., 1982), a poloidal averaging of the standard ballooning equation is carried out to average over the fast variation $L_{\alpha 1}\sim\rho$ and obtain a “distilled ballooning equation”, which involves only the longer length scale $L_{\alpha 2}\sim\hat{s}\rho$. A rigorous multiscale NAE program must be carried out for this purpose, which is beyond the scope of the present work. Here, we shall content ourselves with geometries that depend only weakly on the fast scale, thereby bypassing the need for averaging over the fast scale to obtain the “distilled” equations. Vacuum fields in slab and cylindrical geometries offer such possibilities.
@@
In Section III.1, we first demonstrate the free-surface approach to solving the overdetermination problem of QS in a vacuum field for slab geometry. The analysis is simplified, and lays bare the balance between dispersion and nonlinearity needed to realize soliton solutions (Ablowitz, 2011). Once this structure is understood, it essentially carries over with minimal modification to the more general geometry, which we discuss in III.2.
@@ III.1 Vacuum quasisymmetric magnetic field in slab geometry
where $\hat{q}$ is a small deviation from the low-order rational. The inner solution is trivially quasisymmetric. We will then carry out a surface expansion around the plane $z=h$ with the outermost free-surface parametrized by $z=h+\eta(x,y)$. The amplitude of $\eta(x,y)$ and $h$ will be assumed small compared to the characteristic length $L$ along the field line, which allows us to carry out a multiscale analysis for a weakly nonlinear system typical in the theory of shallow-water waves (Ablowitz, 2011). The justification and the relevance of the assumptions will be made clear in the following.
Exact vacuum fields in slab geometry with closed field lines and nested flux surfaces have been obtained in (Weitzner and Sengupta, 2020). As shown in Weitzner and Sengupta (2020), one of the peculiarities of closed field line systems in a slab geometry is that one can always re-orient the coordinate system such that the magnetic field is purely along one of the axes. For a closed field line, tangential to the $X,Y$ plane and $\iota\equiv(\bm{B}\cdot\bm{\nabla}Y)/(\bm{B}\cdot\bm{\nabla}X)=n/m$, this can be achieved through a combination of rotation and dilation of the coordinates
@@
The system (24) is mathematically similar to some aspects of the theory of shallow-water waves under gravity. Since our problem does not depend explicitly on time, we make $\alpha$ the analog of time. Therefore, in analogy with the shallow-water wave theory (Ablowitz, 2011), we denote the condition (24d) as the “kinematic boundary condition” (KBC), and (24c) as the “generalized Bernoulli” (GB) condition. The differences between water wave theory and QS arise from the difference in the Bernoulli condition and the fact that we treat $y$ (approximately $\alpha$) as a time-like variable.
@@
Guided by the “principle of maximal balance” (Kruskal, 1965; Ablowitz, 2011), we order
@@
In the water-wave theory, the relation (26) represents the delicate balance between nonlinearity and dispersion, which is crucial for the existence of solitons. The significance of (26) in the context of vacuum fields can be clarified by NAE theory. The small parameter $\delta$ is analogous to the near-axis expansion parameter $\kappa\rho$, which is the product of the axis curvature $\kappa$ and distance $\rho$ from the axis. The balance (26) then implies that there are second-order corrections to the NAE flux surface shapes due to the magnetic shear. The difference between the actual surface and the second-order NAE predictions being $O(\delta^{2})$ is consistent with the results obtained using NAE and standard numerical optimization (Landreman, 2022; Rodriguez, 2022). We note that (26) is not the only possible relation for balancing nonlinearity and dispersion (Burde and Sergyeyev, 2013). Other possibilities will be discussed in a future publication.
@@
Here, $Y=\epsilon y$ denotes a longer length scale in $y$. The TW solution is consistent with (6), expanded to the lowest order in $\epsilon$.
@@
Proceeding with the KBC (29d), we find that at the lowest order
@@
Through suitable variable transformations, the above equation can be cast as the standard KdV equation for $B=1+\epsilon\tilde{B}$, i.e.,
@@
Let us point out some of the important predictions of the slab theory. Firstly, the sum $B_{M}+B_{m}+B_{X}$ is determined by the coefficients of KdV (Kamchatnov, 2000), which, as seen from (40) only depends on $q_{0}$. Moreover, as predicted by the NAE theory (Section V) and confirmed numerically (Section VI and Appendix E), the sum of $B_{M},B_{m}$ is essentially a constant equal to the field strength on the axis. Therefore, the third root $B_{X}$, must be determined by the $\iota$ profile. We shall also numerically verify this observation in Section VI.
@@
As shown in Appendix B, a subsidiary expansion leads to the Gardner’s equation (Miura, 1976; Kamchatnov et al., 2012)
@@
where $c_{1}$ is a constant. Due to the additional $B^{2}$ term in the Gardner’s equation (46) compared to the KdV equation (41), under the TW ansatz we need to replace the cubic $f^{2}$ in (42) by a quartic (Kamchatnov et al., 2012). Therefore, we need four roots to describe $B$ in this case, either all real, or two real and a pair of complex conjugates.
@@ III.2 Extension of the slab geometry results to general toroidal geometry
To include the effect of shear, we now go beyond the NAE limit and extend the free-surface idea proposed earlier in the slab model to more general geometries. We seek geometries in which the dependence on the fast variation is straightforward. Such a case arises with a large-aspect-ratio torus with nearly circular flux surfaces. Instead of the Cartesian coordinates, we now use the Clebsch coordinates $(\psi_{s},\alpha_{s},\Phi_{s})$, where the subscript denotes the surface around which we carry out the surface expansion. In these NAE Clebsch coordinates, we have
@@
In summary, the perturbative approach to solving the overdetermined vacuum quasisymmetric system leads to the following results. First, we can show analytically that the compatibility condition between force balance and QS determines the rotational transform profile. Within our approximate model, we are able to show that the compatibility implies that for a given average rotational transform, there is a preference for an optimal shear. In fact, recent work (Buller et al., 2024) has shown that stellarators with precise levels of QS that can be derived from the Landreman-Paul precise QA indeed have an optimal shear for the best QS. We discuss this further in Section VI (see Fig. 5). Second, the functional form of $(\partial_{\ell}B)^{2}$ is not arbitrary. In general, it is cubic in $B$, which leads to the KdV equation for $B$. For small $\iota$ and $N=0$, $(\partial_{\ell}B)^{2}$ is a quartic, leading to Gardner’s equation for $B$.
@@ IV The connection of QS to integrable PDEs: Painlevé property
We note that the first assumption is physically reasonable, has been shown to hold near the axis (Garren and Boozer, 1991a), and is also borne out by extensive computations. The second assumption is an essential requirement for QS. Both of these assumptions do not apply to omnigeneity. Regarding the third assumption, an alternative is to demand that $B$ must satisfy the fundamental TW condition (6) everywhere on the flux surface. The third condition leads to an infinite number of field-line $\alpha$-independent integrals of the form (17). In the perturbative approach, in addition to these assumptions, we also made certain assumptions on the geometry to make analytical progress with the force balance condition.
@@
Before we proceed with our analysis, let us discuss some of the important known exceptional cases of QS. As stated earlier, the case with an exact symmetry (axisymmetric or helically symmetric) is a special case of QS because $\bm{u}$ is determined independently of $\bm{B}$. The third condition, or equivalently $\bm{u}\cdot\bm{\nabla}B=0$, is satisfied identically for every choice of $f(B,\psi)$ in the $\partial_{\ell}B$ equation (10). We find a similar degeneracy in approximate QS near axisymmetry (Sengupta et al., 2024; Plunk and Helander, 2018; Plunk, 2020), where no constraints exist on $\partial_{\ell}B$. The near-axisymmetric devices can not be made quasisymmetric to all orders and have higher symmetry-breaking errors than the Landreman-Paul precise QA. A more general analysis shows that in configurations that are not close to either axisymmetry or helical symmetry, the degeneracy is avoided, and $\partial_{\ell}B$ is once again constrained (Nikulsin et al., 2024). We shall also exclude the exceptional case of zero rotational transform (Sato, 2022) since we focus only on irrational surfaces while assuming continuity on rational surfaces. We shall return to the comparison with the axisymmetric case at the end of this Section.
@@
Our main goal is to study the single-valuedness of $B$ on a constant flux surface $\psi$ that satisfies the $\partial_{\ell}B$ equation (10) and the $\partial_{\alpha}B$ equation (12), subject to the periodicity condition (11). Starting with the pioneering work by S. Kovalevskaya (Conte, 2012) on rigid body dynamics, the study of single-valuedness and singularities of differential equations in the complex domain has yielded fundamental results in several physics problems from quantum field theory to optics (Levi and Winternitz, 2013). Extending from the real to the complex plane is a helpful technique that physicists often utilize in solving linear problems. However, as Kovaleskaya showed, going to the complex plane and analyzing singularities of highly nonlinear systems can be equally insightful. Moreover, singularities in the complex plane, even those that are far from the real line, can determine and control the behavior of analytic functions in the real line. An important example of this is a doubly periodic elliptic function (Hille, 1997; Whittaker and Watson, 1920) for which, singularities determine its periodicity on the real line. Our investigation is primarily motivated by the observation that the hidden symmetry of QS could admit a special Kovalevskaya-like solution (Burby et al., 2020).
@@
To proceed further, we need critical ideas from the well-developed theory of complex ODEs (Ince, 1956; Hille, 1997; Conte, 2012). In contrast to solutions of linear ODEs, whose singularities are determined by those of the coefficients, nonlinear ODEs allow movable singularities, which depend on the initial conditions. The standard example is $y^{\prime}+y^{2}=0$, whose general solution, $y=(x-x_{0})^{-1}$, depends on the integration constant $x_{0}$ determined by the initial condition $y(0)=-x_{0}^{-1}$. The possible singularities of ODEs are known. They include poles, branch points, essential singular points, and essential singular lines. Single-valuedness of functions breaks down near critical singularities, such as branch points of algebraic or logarithmic nature. If the critical singularities are fixed, the functions can be made single-valued by defining suitable branch cuts or fixed Riemann surfaces. However, the presence of singularities, which are both critical and movable, leads to dense multi-valuedness around movable singularities of solutions and nonintegrability (Conte (2012); Kruskal et al. (2007) and references therein). The absence of movable critical singularities in the general solution of an ODE is called the Painlevé property (PP) of an ODE.
We will now look for $\mathcal{B}$ with the PP such that there are no critical singularity in the entire complex plane since their very presence can affect $\mathcal{B}$ on the real line. We now recall some classical results on ODEs with the PP (details given in (Ince, 1956; Hille, 1997; Conte, 2012)). Let us rewrite the $\partial_{\ell}B$ equation in the complex plane as
@@
which is exactly solvable in terms of elliptic functions and their degenerate cases. As is well-known, elliptic functions are doubly-periodic, meromorphic functions. The dependence on the arbitrary initial condition or the TW frame thus leads to only movable poles and no movable critical singularities. Hence, the solution possesses the PP. These classical results follow from standard Painlevé test and extensions of Malmquist’s theorem (Hille, 1997; Eremenko, 1982). They show how stringent and rigid the constraint of single-valuedness can be.
Returning to the QS problem, we first observe that the Riccati equation for $B$ is inconsistent with the NAE solution. The Riccati equation for $B$ will have constant real coefficients on a flux surface, which does not admit nontrivial periodic solutions on the real line Coddington et al. (1956). As discussed in Section III.1 and III.2 the perturbative treatment near the slab or the axis shows that $(\partial_{\ell}B)^{2}=P_{3}(B),P_{4}(B)$ is more appropriate, which coincides with two of the four forms of ODEs of first order and second degree with PP. The periodicity condition can be easily satisfied with the elliptic functions for these cases. Now, the cubic and quartic polynomials are the twice-integrated forms of the KdV and Gardner’s equations. The PP of the KdV and Gardner’s equation in the TW frame is well-known and is conjectured to be essential for integrable PDEs with soliton solutions (Ablowitz et al., 1980a, b).
We now address whether PP is necessary and sufficient for a quasisymmetric $B$. It is certainly sufficient for QS since both the triple product form (8) and the periodicity condition can be satisfied. However, the necessity is not apparent, particularly in light of axisymmetry, where no such stringent restriction on $\partial_{\ell}B$ appears. In axisymmetry, weak singularities such as $R^{m}\log{R}^{n}$, with integer $m,n$ often arise in the description of $\psi$ (Cerfon and Freidberg, 2010) and in $B$ in cylindrical coordinates. However, these singularities are fixed and, therefore, not of the critical movable type. To further understand the difference between a perfectly axisymmetric system and a system with excellent but not exact QS, we need to look deeper into the $\bm{u}\cdot\bm{\nabla}B=0$ condition. As discussed in Appendix A, if QS were exact, we would have the same characteristics as axisymmetry. However, unlike the axisymmetric problem, the QS problem is intrinsically overdetermined and three-dimensional. With this in mind, we now motivate the PP as a criterion that allows one to look for robust quasisymmetric systems.
@@
Figure 1: The value of $(dB/d\ell)^{2}$ on the outermost flux surface in the precise QA equilibrium Landreman and Paul (2022) is plotted as a function of $B$, alongside a cubic polynomial fit in (a). In (b), the errors of quadratic, cubic, and quartic fits are compared on each flux surface, demonstrating that a cubic ansatz (as predicted by KdV) is sufficient. Hence, precise QA has the Painlevé property for $B$.
Figure 3: Prediction of maximum toroidal volume that can be made quasisymmetric for precise QA. The boundary of the extended precise QA and two other flux surfaces are shown inside a Poincare plot of the islands in (a), and the roots are plotted in (b). The boundary of the original precise QA is shown in red. Note that the islands occur where the second and third roots touch approximately.